# Decentralized Federated Learning: Model Update Tracking Under Imperfect Information Sharing

### 20 Mar 2024 v1

@@ III Convergence Analysis
$\hskip 11.38109pt+\frac{\rho}{8n}\mathbb{E}||X^{t}-\bar{X}^{t}||_{F}^{2}+\frac{\eta^{2}\rho\sigma^{2}(1-\mu)}{8}+\frac{\eta^{2}\rho D^{t,2}}{8}+\frac{\eta^{2}\rho\zeta^{2}}{8}$
Theorem 1 (Smooth non-convex cases for FedNMUT).
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}||\nabla f(\bar{x}^{t})||^{2}\leq\frac{2}{\eta T}(f(\bar{x}^{0}-f^{*})$
$\hskip 7.11317pt+2L^{2}\eta^{2}\sigma^{2}\Big{[}\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\Big{]}+2L^{2}\eta^{2}\zeta^{2}\Big{[}\frac{48}{\rho^{2}}+\frac{1}{2}\Big{]}$
$\hskip 7.11317pt+2L^{2}\eta^{2}D^{t,2}\Big{[}\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\Big{]}$
where all expectations are w.r.t. the data and the noise.
Corollary 1.1.
@@
Suppose that the step size satisfies $\eta=\mathcal{O}\Big{(}\sqrt{\frac{n}{T}}\Big{)}$ For a sufficiently large $T$ we have,
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}||\nabla f(\bar{x}^{t})||^{2}\leq\mathcal{O}\Bigg{(}\frac{1}{\sqrt{nT}}+\frac{1}{T}\Bigg{)}.$ (8)
Theorem 1 establishes a worst-case upper bound on the convergence of the algorithm studied in the paper. In particular, the theorem jointly bounds the expected gradient norm, which is a notion of approximate first-order stationarity of the average iterate $\bar{x}_{t}$, and the consensus error. The convergence bounds consist of four terms: the first term effectively captures the error arising from inaccurate initialization. The second and third term captures the effect of stochasticitiy and data heterogeneity. The last term captures the adverse effect of imperfect communication modeled as communication noise.
The finding presented in Corollary 1.1 shows that the FedNMUT algorithm can achieve a linear speedup, exhibiting a convergence rate of $\mathcal{O}(\frac{1}{\sqrt{nT}})$, given that $T$ is large enough and unaffected by the communication topology. This rate of convergence is comparable to the best-known results for decentralized SGD algorithms found in existing literature.
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 5: Loss vs. iterations with and without noise (Var=0.01) for different communication topologies. $\mu$ = 0.02
(a) Fully-connected topology
(b) Torus topology
(c) Ring topology
Figure 6: Consensus error vs. iterations with and without noise (Var=0.01) for different communication topologies. $\mu$ = 0.02. Note the different Y-axis scale in Fig (a) as compared with Fig (b) and (c) for better readability.
IV Experiments
We conduct a series of experiments on regression tasks to evaluate the effects of noise on the convergence behavior of the algorithm outlined in Theorem 1. We set the number of clients, denoted as $n$, to 16. Each experiment is executed thrice, with the outcomes (loss/consensus error) being averaged. Our approach employs a mean-squared error loss function enhanced with $L_{2}$ regularization. We establish the initial learning rate at 0.2, applying a decay factor of 0.9 in subsequent iterations. We create data samples ($m$ = 10,000) in the form of ${(x_{i};y_{i})}^{m}_{i=1}$, modeled as $y_{i}=\langle w,x_{i}\rangle+\epsilon_{i}$, where $w$ belongs to $\mathbb{R}^{2000}$, $x_{i}$ follows a normal distribution $\mathcal{N}(0;I_{2000})$, and noise $\epsilon_{i}$ adheres to $\mathcal{N}(0,0.05)$.
We commence with noise-free trials to establish a baseline, subsequently incrementing noise variance to examine algorithmic robustness. For consistency, the experimental findings with noise variances of $D^{2,t}=0.005$ and $D^{2,t}=0.01$ are depicted alongside the no-noise condition in Figures 3, 4, 5 and 6. These tests were conducted using an Intel Xeon Gold workstation.
The consensus error depends on the topology of the communication network. We observed that the consensus error is low for the fully connected network and high for the ring topology for the same algorithm in the presence of noise which is also consistent with the Theorem 1.
V Conclusion
We studied the impact of noisy communication channels on the convergence of Decentralized Federated Learning with Model Update Tracking method. We proposed multiple scenarios for establishing consensus in the presence of noise and provided experimental results on the algorithm. Additionally, we provided theoretical results for FedNMUT under the assumption of smooth non-convex function, and we observed that the noise term in the upper bound given by Theorem 1 is of order $\mathcal{O}(\frac{1}{\sqrt{nT}})$ and is independent of communication topology. We conducted numerical experiments and observed that FedNMUT is more robust against the added noise than the existing state-of-the-art methods that include communication noise.

### 24 Mar 2024 v2

@@ Appendix C Main Results
$\begin{split}&\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\leq\frac{2}{\eta T}(f(\bar{x}^{0}-f^{*})+\frac{L\mu^{2}\eta}{nT}\sum_{t=0}^{T-1}\mathbb{E}\|B^{t}\|_{F}^{2}+2L^{2}\eta^{2}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\right]\\ &\hskip 113.81102pt+2L^{2}\eta^{2}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]+\frac{2L^{2}\eta^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=1}^{n}{D}^{2}_{t,i}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\right]\end{split}$ (12)
where $f(\bar{x}^{0})-f^{*}$ is the sub-optimality gap, $\bar{x}$ is the average/consensus model parameters.
Proof:
Recall Lemma 3
$\frac{1}{n}\mathbb{E}\|X^{t+1}-\bar{X}^{t+1}\|_{F}^{2}\leq$ $\frac{1-\rho/4}{n}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+{4\eta^{2}\sigma^{2}}+\frac{12\eta^{2}\zeta^{2}}{\rho}+\frac{6\eta^{2}\mu^{2}}{n\rho}\mathbb{E}\|B^{t}\|_{F}^{2}$
$+\frac{4\eta^{2}}{n\rho}\sum_{i=1}^{n}{D}^{2}_{t,i}.$
and Lemma 4
$\frac{6\eta^{2}\mu^{2}}{n\rho(1-\mu)}\mathbb{E}\|B^{t+1}\|_{F}^{2}$ $\leq\Big{(}\frac{6\eta^{2}\mu^{2}}{n\rho(1-\mu)}-\frac{6\eta^{2}\mu^{2}}{n\rho}\Big{)}\mathbb{E}\|B^{t}\|_{F}^{2}+\frac{\rho}{8n}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}$
$+\frac{\eta^{2}\rho\sigma^{2}(1-\mu)}{8}+\frac{\eta^{2}\rho}{8n}\sum_{i=1}^{n}{D}^{2}_{t,i}+\frac{\eta^{2}\rho\zeta^{2}}{8}.$
Combining Lemma 3 and Lemma 4 and simplifying, we obtain
$\frac{1}{n}\mathbb{E}\|X^{t+1}-\bar{X}^{t+1}\|_{F}^{2}+\frac{6\eta^{2}\mu^{2}}{n\rho(1-\mu)}\mathbb{E}\|B^{t+1}\|_{F}^{2}\leq$ $\frac{1-\rho/4}{n}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+{4\eta^{2}\sigma^{2}}+\frac{12\eta^{2}\zeta^{2}}{\rho}$
$+\frac{6\eta^{2}\mu^{2}}{n\rho}\mathbb{E}\|B^{t}\|_{F}^{2}+\frac{4\eta^{2}}{n\rho}\sum_{i=1}^{n}{D}^{2}_{t,i}$
$+\Big{(}\frac{6\eta^{2}\mu^{2}}{n\rho(1-\mu)}-\frac{6\eta^{2}\mu^{2}}{n\rho}\Big{)}\mathbb{E}\|B^{t}\|_{F}^{2}+\frac{\rho}{8n}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}$
$+\frac{\eta^{2}\rho\sigma^{2}(1-\mu)}{8}+\frac{\eta^{2}\rho}{8n}\sum_{i=1}^{n}{D}^{2}_{t,i}+\frac{\eta^{2}\rho\zeta^{2}}{8}$
Simplifying and multiplying the above equation by $\frac{4L^{2}\eta}{\rho}$ gives
$\frac{4L^{2}\eta}{n\rho}\mathbb{E}\|X^{t+1}-\bar{X}^{t+1}\|_{F}^{2}+\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}(1-\mu)}\mathbb{E}\|B^{t+1}\|_{F}^{2}$
$\leq\left[\frac{4L^{2}\eta}{\rho n}-\frac{L^{2}\eta}{n}\right]\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{16L^{2}\eta^{3}\sigma^{2}}{\rho n}+\frac{48L^{2}\eta^{3}\zeta^{2}}{\rho^{2}}+\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}}\mathbb{E}\|B^{t}\|_{F}^{2}+\frac{16L^{2}\eta^{3}}{n\rho^{2}}D^{2,t}$
$+\Big{(}\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}(1-\mu)}-\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}}\Big{)}\mathbb{E}\|B^{t}\|_{F}^{2}+\frac{L^{2}\eta}{2n}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{L^{2}\eta^{3}\sigma^{2}(1-\mu)}{2}+\frac{L^{2}\eta^{3}D^{t,2}}{2}+\frac{L^{2}\eta^{3}\zeta^{2}}{2}$
$\leq\left[\frac{4L^{2}\eta}{n\rho}-\frac{L^{2}\eta}{2n}\right]\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}(1-\mu)}\mathbb{E}\|B^{t}\|_{F}^{2}+L^{2}\eta^{3}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}\right]+L^{2}\eta^{3}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]$
$\hskip 56.9055pt+L^{2}\eta^{3}\sum_{i=1}^{n}{D}^{2}_{t,i}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}\right]$
Let
$\begin{split}\Phi^{t}=\frac{4L^{2}\eta}{n\rho}\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}(1-\mu)}\mathbb{E}\|B^{t}\|_{F}^{2}+\mathbb{E}[f(\bar{x}^{t})-f^{*}]\end{split}$ (13)
Recall now Lemma 2,
$\mathbb{E}f(\bar{x}^{t+1})$ $\leq\mathbb{E}f(\bar{x}^{t})+\frac{L\eta^{2}\sigma^{2}}{2n}-\frac{3\eta}{8}\ \mathbb{E}\|\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{t})\|^{2}-\frac{\eta}{2}\ \mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}$
$\hskip 56.9055pt+\frac{L^{2}\eta}{2n}\ \mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{L\eta^{2}}{2n}D^{2,t}+\frac{L\mu^{2}\eta^{2}}{2n}\mathbb{E}\|B^{t}\|_{F}^{2}$
Combining Lemma 2 and Equation 13, we have the following
$\begin{split}\Phi^{t+1}&\leq\left[\frac{4L^{2}\eta}{n\rho}-\frac{L^{2}\eta}{2n}\right]\mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{24L^{2}\eta^{3}\mu^{2}}{n\rho^{2}(1-\mu)}\mathbb{E}\|B^{t}\|_{F}^{2}\\ &\hskip 56.9055pt+L^{2}\eta^{3}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}\right]+L^{2}\eta^{3}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]+L^{2}\eta^{3}D^{t,2}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}\right]\\ &\hskip 56.9055pt+\frac{L\eta^{2}\sigma^{2}}{2n}-\frac{3\eta}{8}\ \mathbb{E}\|\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{t})\|^{2}-\frac{\eta}{2}\ \mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\\ &\hskip 56.9055pt+\frac{L^{2}\eta}{2n}\ \mathbb{E}\|X^{t}-\bar{X}^{t}\|_{F}^{2}+\frac{L\eta^{2}}{2n}D^{2,t}+\frac{L\mu^{2}\eta^{2}}{2n}\mathbb{E}\|B^{t}\|_{F}^{2}\\ &\leq\Phi^{t}+L^{2}\eta^{3}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\right]+L^{2}\eta^{3}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]\\ &\hskip 28.45274pt+L^{2}\eta^{3}D^{t,2}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\right]+\frac{L\mu^{2}\eta^{2}}{2n}\mathbb{E}\|B^{t}\|_{F}^{2}\\ &\hskip 28.45274pt-\frac{3\eta}{8}\ \mathbb{E}\|\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{t})\|^{2}-\frac{\eta}{2}\ \mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\\ \implies&\frac{\eta}{2}\mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\leq(\Phi^{t}-\Phi^{t+1})+\frac{L\mu^{2}\eta^{2}}{2n}\mathbb{E}\|B^{t}\|_{F}^{2}+L^{2}\eta^{3}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\right]+\\ &\hskip 14.22636ptL^{2}\eta^{3}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]+L^{2}\eta^{3}D^{t,2}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\right]-\frac{3\eta}{8}\ \mathbb{E}\|\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{t})\|^{2}\end{split}$
Telescoping over iterations $t$ from 00 to $T$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\leq\frac{2}{\eta T}(f(\bar{x}^{0}-f^{*})+\frac{L\mu^{2}\eta}{nT}\sum_{t=0}^{T-1}\mathbb{E}\|B^{t}\|_{F}^{2}+2L^{2}\eta^{2}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\right]$
$\hskip 113.81102pt+2L^{2}\eta^{2}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]+\frac{2L^{2}\eta^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=1}^{n}{D}^{2}_{t,i}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\right]$ (14)
This concludes the proof of the Theorem 2.
C-A Proof of Corollary
@@
Suppose that the step size $\eta=\mathcal{O}\Big{(}\sqrt{\frac{n}{T}}\Big{)}$, then for a sufficiently large $T$, we have $\bar{B}^{2}=\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|B^{t}\|_{F}^{2}$ and $\bar{D}^{2}=\frac{1}{nT}\sum_{t,i=1,1}^{T,n}D^{2}_{t,i}$
$\begin{split}(i)\hskip 5.69054pt&\eta\leq\min\Big{\{}\frac{1}{4L},\frac{\rho}{7L}\Big{\}}\\ (ii)\hskip 5.69054pt&\frac{\mu}{1-\mu}\leq\frac{\rho}{42}\\ (iii)\hskip 5.69054pt&\frac{6\mu^{2}}{\rho(1-\mu)}\leq\frac{\rho}{8}\end{split}$
If the step size $\eta$ is $\mathcal{O}(\sqrt{\frac{n}{T}})$, then we have the following order of convergence for each term in Theorem 1:
$\frac{2}{\eta T}(f(\bar{x}^{0}-f^{*})=\mathcal{O}\left[\frac{1}{\sqrt{nT}}\right]=\mathcal{O}\left[\frac{1}{\sqrt{T}}\right]$
$\frac{L\mu^{2}\eta}{nT}\sum_{t=0}^{T-1}\mathbb{E}\|B^{t}\|_{F}^{2}$ $=\mathcal{O}\left[\frac{1}{\sqrt{nT}}\right]\bar{B}^{2}=\mathcal{O}\left[\frac{1}{\sqrt{T}}\bar{B}^{2}\right]$
$2L^{2}\eta^{2}\sigma^{2}\left[\frac{16}{\rho n}+\frac{1-\mu}{2}+\frac{1}{2nL\eta}\right]$ $=\frac{2L^{2}\eta^{2}\sigma^{2}}{\rho n}+L^{2}\eta^{2}\sigma^{2}(1-\mu)+\frac{L\eta\sigma^{2}}{n}$
$=\mathcal{O}\left[\frac{1}{\sqrt{T}}\sigma^{2}\right]$
$2L^{2}\eta^{2}\zeta^{2}\left[\frac{48}{\rho^{2}}+\frac{1}{2}\right]$ $=\frac{96L^{2}\eta^{2}\zeta^{2}}{\rho^{2}}+L^{2}\eta^{2}\zeta^{2}$
$=\mathcal{O}\left[\frac{1}{T}\zeta^{2}\right]$
$\frac{2L^{2}\eta^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=1}^{n}{D}^{2}_{t,i}\left[\frac{16}{n\rho^{2}}+\frac{1}{2}+\frac{1}{2nL\eta}\right]$ $=\frac{32L^{2}\eta^{2}\bar{D}^{2}nT}{n\rho^{2}T}+\frac{L^{2}\eta^{2}\bar{D}^{2}nT}{T}+\frac{L\eta\bar{D}^{2}nT}{nT}$
$=\mathcal{O}\left[\frac{1}{\sqrt{T}}\bar{D}^{2}\right]$
The overall convergence rate is
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f(\bar{x}^{t})\|^{2}\leq\mathcal{O}\left[\frac{1}{\sqrt{T}}\bar{B}^{2}+\frac{1}{\sqrt{T}}\sigma^{2}+\frac{1}{T}\zeta^{2}+\frac{1}{\sqrt{T}}\bar{D}^{2}\right],$
Therefore, at large $T$, the convergence rate of FedNMUT is $\mathcal{O}(\frac{1}{\sqrt{T}})$.